Given that real numbers x,y,z satisfy the system of equations
x+y+z=6
x^2+y^2+z^2=26
x^3+y^3+z^3=90
Find the values of x×y×z and x^4+y^4+z^4
1
Expert's answer
2019-12-31T07:00:54-0500

"36=(x+y+z)^2=(x^2+y^2+z^2)+2(xy+yz+zx)=26+2(xy+y z+zx)"

"\u21d2xy+yz+zx=(36\u221226)\/2=10\/2=5"


Using formula

"(x+y+z)(x^2+y^2+z^2-xy\u2212yz\u2212zx)=x^3+y^3+z^3\u22123xyz"

We can find xyz:

"6(26\u22125)=90\u22123xyz,126=90\u22123xyz,xyz=(90\u2212126)\/3=\u221236\/3=\u221212"


Using formula "x^4+y^4+z^4=(x+y+z)^4\u22124(x+y+z)^2(xy+yz+zx)\n+2(xy+yz+zx) ^2\n +4(x+y+z)xyz"


We can find "x ^4 +y ^4 +z \n^4 =6 ^4 \u22124\u22176 ^2 \u22175+2\u22175 ^2\n +4\u22176\u2217(\u221212)=338"

Answer: "xyz=\u221212,x \n^4 +y ^4 +z ^4 =338"